calculate the pH of a 1M weak acid solution for which pKa=6
Answer #1

Setting up an iCE table of weak acid HA with fractional dissociation x, we get:

                                            HA    \rightarrow          H+     +      A-

Initial concentration:              1                   0               0

Equilibrium concentration:   1 - x                x                x

We have the relation: pKa = -log(Ka)

Ka = 10-pKa = 10-6

Writing the equation for Ka, Ka = [H+} * [A-] / [HA]

Using the values from the table: 10-6 = x * x / (1-x)

Solving, x = [H+] = 10-3

pH = -log( [H+]) = -log(10-3) = 3